Hi my son is having problems with this worksheet it says fractions and decimals on it but directions say'set-up and solve using Let. Statements and Eq. Statements solve by elemenating the fractions or the decimals. Write a word phrase and I can't seem to figure it out can u help us on the first on please.
The sum of 3.8 and twice a number is 12.4 . What's the number?
Can u show us how u did it for the other problems he has thank you Mrs.Bartlett

Answer by Edwin McCravy(20054)

The sum of 3.8 and twice a number is 12.4
Substitute the letter x for the words "a number"
The sum of 3.8 and twice x is 12.4
Substitute 2x for the words "twice x"
The sum of 3.8 and 2x is 12.4
Substitute an equal sign "=" for the word "is"
The sum of 3.8 and 2x = 12.4
Substitute 3.8+2x for the words "The sum of 3.8 and 2x"
3.8+2x = 12.4
Remove the decimals by multiplying every term by 10

(10)(3.8)+(10)(2x) = (10)(12.4)
38 + 20x = 124
Subtract 38 from both sides

20x = 86
x = 86/20
x = 43/10
x = 4.3
Edwin